OpenStudy (anonymous):

3x^(5/3)=0 will fan and medal!!!

OpenStudy (anonymous):

Simplifying x + 2 = 3x + -1x2 + 5 Reorder the terms: 2 + x = 3x + -1x2 + 5 Reorder the terms: 2 + x = 5 + 3x + -1x2 Solving 2 + x = 5 + 3x + -1x2 Solving for variable 'x'. Reorder the terms: 2 + -5 + x + -3x + x2 = 5 + 3x + -1x2 + -5 + -3x + x2 Combine like terms: 2 + -5 = -3 -3 + x + -3x + x2 = 5 + 3x + -1x2 + -5 + -3x + x2 Combine like terms: x + -3x = -2x -3 + -2x + x2 = 5 + 3x + -1x2 + -5 + -3x + x2 Reorder the terms: -3 + -2x + x2 = 5 + -5 + 3x + -3x + -1x2 + x2 Combine like terms: 5 + -5 = 0 -3 + -2x + x2 = 0 + 3x + -3x + -1x2 + x2 -3 + -2x + x2 = 3x + -3x + -1x2 + x2 Combine like terms: 3x + -3x = 0 -3 + -2x + x2 = 0 + -1x2 + x2 -3 + -2x + x2 = -1x2 + x2 Combine like terms: -1x2 + x2 = 0 -3 + -2x + x2 = 0 Factor a trinomial. (-1 + -1x)(3 + -1x) = 0 Subproblem 1 Set the factor '(-1 + -1x)' equal to zero and attempt to solve: Simplifying -1 + -1x = 0 Solving -1 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '1' to each side of the equation. -1 + 1 + -1x = 0 + 1 Combine like terms: -1 + 1 = 0 0 + -1x = 0 + 1 -1x = 0 + 1 Combine like terms: 0 + 1 = 1 -1x = 1 Divide each side by '-1'. x = -1 Simplifying x = -1 Subproblem 2 Set the factor '(3 + -1x)' equal to zero and attempt to solve: Simplifying 3 + -1x = 0 Solving 3 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + -1x = 0 + -3 Combine like terms: 3 + -3 = 0 0 + -1x = 0 + -3 -1x = 0 + -3 Combine like terms: 0 + -3 = -3 -1x = -3 Divide each side by '-1'. x = 3 Simplifying x = 3 Solution x = {-1, 3}